[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Oct 09, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Oct 9 08:37:26 UTC 2017


HSU Shift Log for Mon Oct 09 2017 04:37:17 GMT-0400 (EDT)

Shift: A

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues: Not much of an issue, but just wanted to pass on that the aids  
colors are not the standard ones.

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171009/613b6496/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list